OpenStudy (anonymous):

simplify by taking roots of numerator and denominatior: √100/121 = my answer is 10/11 is this correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnks satellite73 where have you been good to see u on here

OpenStudy (anonymous):

satellite73 check this for me multiply and simply by factoring ^4√10,000 * ^4√100,000 = 160000√10

OpenStudy (anonymous):

vacation

OpenStudy (anonymous):

is this \[\sqrt[4]{10^4}\times \sqrt[4]{10^5}\]?

OpenStudy (anonymous):

if so, answer is \[\sqrt[4]{10^9}\] giving \[100\sqrt[4]{10}\]

OpenStudy (anonymous):

no ^4√10,000 * ^4√100,000 =

OpenStudy (anonymous):

does this mean fourth root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

excuse me i meant \[10000=10^4\]

OpenStudy (anonymous):

and likewise \[100000=10^5\]

OpenStudy (anonymous):

and further \[10^4\times 10^5=10^9\] all this should be more or less clear i hope

OpenStudy (anonymous):

and you want the fourth root of this big number. we reason as follows: 4 goes in to 9 two times with a remainder of 1. so two tens come out of the radical, one stays in gives \[10^2\times \sqrt[4]{10}\]

OpenStudy (anonymous):

just like \[\sqrt[4]{x^9}=x^2\sqrt[4]{x}\]

OpenStudy (anonymous):

so the ans. is 100 ^4√10

OpenStudy (anonymous):

yes for sure